# Schrödinger Symmetry in Spherically-symmetric Static Mini-superspaces with Matter Fields

### v1

@@ VI.2 $\{G,H\}=0$: Solution transformations
We use the mini-superspace models discussed so far and demonstrate that generators $G$ with $\{G,H\}=0$ map a physical solution to another in the same theory with $H$, where both satisfy $H\approx 0$. Among the generators of the Schrödinger symmetry, $\widetilde{P}_{a}$ and $\widetilde{J}_{ab}$ are of this type. We will pick up particularly interesting cases.
@@
and study the transformation of the (A)dS Reissner-Nordström solution (IV.2) by $\widetilde{J}_{-A}$. In particular, let us map this to the (A)dS Schwarzschild solution (III.3) with a different value of mass $M$.
@@
$=-\frac{2}{\epsilon cl_{p}}\left(\alpha_{+}+\frac{(-\epsilon)\alpha_{A}^{2}}{4\alpha_{-}}\right)\alpha_{+}-\frac{cl_{p}}{2l_{s}^{2}}$
@@
Therefore, we reach the AdS-Schwarzschild solution with mass $M^{\prime}$. This shows that the symmetry transformation by $J_{A-}$ for $s=s_{*}$ maps the AdS-RN solution with $(M,\mathcal{Q})$ to the AdS-Schwarzschild solution with $M^{\prime}$ in the same Hamiltonian constraint $H_{gE}\approx 0$.
@@
where in the last step we have used $H_{g\phi}|_{\widetilde{\Pi}_{a}\to\alpha_{a}}=0$. Therefore, the transformed configuration $Y^{\prime a}(\tau)=\beta^{\prime a}+\mathcal{G}^{ab}_{\phi}\alpha^{\prime}_{b}\tau$ satisfies the original Hamiltonian constraint. Here, the $\mu$ parameter (V.3) changes as $\mu=\frac{\alpha_{0}}{\alpha_{1}}>1\to\frac{\alpha^{\prime}_{0}}{\alpha^{\prime}_{1}}=1$ because of $\alpha^{\prime}_{\phi}=0$. Therefore, this is the Schwarzschild solution with a different value of mass (111), $M^{\prime}=\frac{l_{s}\alpha^{\prime}_{1}}{4G\bar{c}}e^{-2\left(\beta^{\prime 0}+\beta^{\prime 1}\right)}$.
@@ VI.3 $\{G,H\}\neq 0$: Theory transformations
We first note from footnote 6 that the original relations $\widetilde{X}^{a}(X,P),\widetilde{P}_{a}(X,P)$ are kept in the new theory, and the same geometric interpretation holds except for the parts involving the Hamiltonian constraint. Then, the calculation for $g_{rr}$ (IV.2) is modified as
@@
to the Einstein-Hilbert action coupled with single massless scalar field $\phi_{I=1}=\phi$ (V.1) where $\Lambda=0$, $m_{I}=0$, and $I=1$. This is the new theory. In other words, this can be regarded as an effective theory obtained by introducing another field $\psi$ with the same excitation as the excitation of field $\phi$ in the original configuration and then integrating it out. The new configuration with the JNW metric discussed above is a solution of this new theory, due to the background field $\psi_{\rm back}$ which replaces the effect of the original field $\phi$. Better understanding the physical meaning of this would require further research.
@@ VII Conclusions and Prospects
Symmetry should be a useful concept in exploring the principles of quantum gravity. The gauge symmetry of spacetime is general covariance, and Hamiltonian constraints play a central role in the dynamics of gravity. Then, what is the physical symmetry in the configuration space of gravity? Hamiltonian constraints define an infinite-dimensional configuration space, “superspace”. Imposing geometric symmetries (such as homogeneity) on spacetime yields a finite-dimensional configuration space, “mini-superspace”. This reduction procedure can be viewed as a kind of fluid limit for (quantum) gravitational configurations. Generically, when many microscopic degrees of freedom are coarse-grained into a fluid description, symmetries like volume-preserving diffeomorphism invariance often emerge, independent of the details of the constituents. Therefore, it is expected that some common symmetry may emerge in various mini-superspaces, potentially carrying the universal symmetry of quantum gravity. Indeed, 2D Schrödinger symmetry was obtained in the homogeneous isotropic universe model and the spherically symmetric static vacuum model. In this paper, to investigate its robustness, we analyzed spherically symmetric static mini-superspace models coupled with matter fields at a classical level. Due to spherical symmetry and staticity, these become one-dimensional dynamical systems with Hamiltonian constraint $H=N\mathcal{H}\approx 0$, where the radial coordinate is the evolution parameter.
@@
First, it is crucial to investigate the robustness of Schrödinger symmetry in a broader context. Our current discussion is limited to cases reducible to the free-particle Hamiltonian without potential via canonical transformations. However, potentials of special forms, such as $V(x)=\frac{k}{x^{2}}$, also preserve Schrödinger symmetry Boyer:1974qt ; Nikitin:2017gtu ; Duval:2024eod . Therefore, it is important to consider such more general cases and investigate whether other matter fields, such as massive scalar field or Dirac field, possess Schrödinger symmetry, or reconsider whether the scalar-field model in terms of the $X$ coordinate in footnote 4 leads to the symmetry. Conversely, it would be interesting to assume such special potentials and examine the spacetime structure derived from them or consider the corresponding modified gravity theories.
@@
In particular, quantization of the scalar-field mini-superspace model is interesting. The interior region of the horizon in the Schwarzschild metric corresponds—via swapping two coordinates $t$ and rr— to a Kantowski-Sachs-type cosmological spacetime, where no matter field exists, of course. From this corresponding, quantization models of black holes are expected to exhibit a strong link with quantum cosmology models Modesto:2005zm . On the other hand, a scalar field can be used as a “clock” in relational formalism to quantize models Domagala:2010bm ; Ashtekar:2011ni . One of our interesting results in the scalar-field mini-superspace model is that, the JNW solution and the Kantowski-Sachs-type spacetime (with the causal-limit fluid) are unified into the single metric (V.3), meaning that the latter corresponds to the “interior” of the former. In particular, both possess the Schrödinger symmetry. Therefore, the quantization of this model based on the symmetry and this interior-exterior relationship should provide a promising avenue for connecting quantum black hole models and quantum universe models.
@@
Finally, it is crucial to examine the Schrödinger symmetry found in our mini-superspace models from the perspective of a hydrodynamic approximation derived from a full theory of quantum gravity. Ref.Oriti:2024elx focuses on the Schrödinger symmetry in the FLRW mini-superspace model and discusses the potential for removing cosmological singularities and realizing accelerated expansion through corrections to the quantum fluid equations that preserve the symmetry. It would be interesting to check whether a similar mechanism works in the spherically-symmetric static mini-superspace models. This test could serve as a crucial benchmark for assessing the universality of the hydrodynamic picture in quantum gravity. On the other hand, Ref.BenAchour:2023dgj pointed out the existence of Schrödinger symmetry in the spherically-symmetric static vacuum mini-superspace and suggested that nonlinear terms preserving the symmetry could serve as quantum corrections. However, the form of such corrections may depend on the choice of the lapse function. Indeed, the symmetry they identified corresponds to only one of the two possible choices of the lapse function and mini-superspace coordinate we found. It would be intriguing to consider different lapses and coordinates and investigate whether there are differences in the physical effects arising from the correction terms. Furthermore, we showed that Schrödinger symmetry persists in spherically-symmetric static spacetimes coupled with matter fields. If the emergence of Schrödinger symmetry in such a mini-superspace is rigorously linked to the validity of the hydrodynamic approximation in a full quantum gravity theory incorporating matter fields, the findings of this study would provide significant implications.
@@ Acknowledgments
We would like to thank Jibril Ben Achour, Che-Yu Chen, Kristina Giesel, Philipp Höhn, Christy Kelly, Etera Livine, Francesco Sartini, and Shinji Tsujikawa for fruitful and inspiring discussions. T.S. is supported by Waseda University Special Research Projects (Nos. 2025C-493) and the 15th Early Bird Program of Waseda University. Y.Y. appreciated the wonderful hospitality of the members of the Institute for Theoretical Physics III at the Friedrich-Alexander-University Erlangen-Nuremberg, where he had a fruitful discussion and completed this work. Y.Y. was partially supported by Japan Society for the Promotion of Science (No.21K13929).
@@ Appendix A Mass and Electric Charge
where $g_{2d}$ is the determinant of the 2D metric normal to 2-sphere, and $dt_{\infty}=kdt$ is the time coordinate at $r=\infty$ with a constant $k$.
@@
Next, we show that in the Y-coordinate, the Komar formula (153) becomes the simpler formula (111). In terms of $dt_{\infty}=\frac{l_{s}\widetilde{\Pi}_{1}}{\bar{c}r_{0}}dt\equiv\frac{1}{k}dt$ and the gauge (85), the Y-metric (81) is expressed as

### v2

@@ VI.2 $\{G,H\}=0$: Solution transformations
We use the mini-superspace models discussed so far and demonstrate that generators $G$ with $\{G,H\}=0$ map a physical solution to another in the same theory with $H$, where both satisfy $H\approx 0$. Among the generators of Schrödinger symmetry, $\widetilde{P}_{a}$ and $\widetilde{J}_{ab}$ are of this type. We will pick up particularly interesting cases.
@@
and study transformation of a (A)dS-RN solution (IV.2) by $\widetilde{J}_{-A}$. In particular, let us map it to a (A)dS Schwarzschild solution (III.3) with a different value of $M$.
@@
$=-\frac{2}{\epsilon cl_{p}}\left(\alpha_{+}+\frac{(-\epsilon)\alpha_{A}^{2}}{4\alpha_{-}}\right)\alpha_{-}-\frac{cl_{p}}{2l_{s}^{2}}$
@@
Therefore, we reach the AdS-Schwarzschild solution with mass $M^{\prime}$. This shows that the symmetry transformation by $\widetilde{J}_{A-}$ for $s=s_{*}$ maps the AdS-RN solution with $(M,\mathcal{Q})$ to the AdS-Schwarzschild solution with $M^{\prime}$ in the same Hamiltonian constraint $H_{gE}\approx 0$.
@@
where in the last step we have used $H_{g\phi}|_{\widetilde{\Pi}_{a}\to\alpha_{a}}=0$. Therefore, the transformed configuration $Y^{\prime a}(\tau)=\beta^{\prime a}+\mathcal{G}^{ab}_{\phi}\alpha^{\prime}_{b}\tau$ satisfies the original Hamiltonian constraint. Here, the $\mu$ parameter (100) changes as $\mu=\frac{\alpha_{0}}{\alpha_{1}}>1\to\frac{\alpha^{\prime}_{0}}{\alpha^{\prime}_{1}}=1$ because of $\alpha^{\prime}_{\phi}=0$. Therefore, this is the Schwarzschild solution with a different value of mass (112), $M^{\prime}=\frac{l_{s}\alpha^{\prime}_{1}}{4G\bar{c}}e^{-2\left(\beta^{\prime 0}+\beta^{\prime 1}\right)}$.
@@ VI.3 $\{G,H\}\neq 0$: Theory transformations
We first note from footnote 7 that the original relations $\widetilde{X}^{a}(X,P),\widetilde{P}_{a}(X,P)$ are kept in the new theory, and the same geometric interpretation holds except for the parts involving the Hamiltonian constraint. Then, the calculation for $g_{rr}$ (IV.2) is modified as
@@
to the Einstein-Hilbert action coupled with single massless scalar field $\phi_{I=1}=\phi$ (V.1) where $\Lambda=0$, $m_{I}=0$, and $I=1$. This is the new theory. In other words, this can be regarded as an effective theory obtained by introducing another field $\psi$ with the same excitation as the excitation of field $\phi$ in the original configuration and then integrating it out. The new configuration with the JNW metric is a solution of this new theory, due to the background field $\psi_{\rm back}$ which replaces the effect of the original field $\phi$. Better understanding the physical meaning of this would require further research.
@@ VII Conclusions and Prospects
Symmetry should be a useful concept in exploring the principles of quantum gravity. The gauge symmetry of spacetime is general covariance, and Hamiltonian constraints play a central role in the dynamics of gravity. Then, what is the physical symmetry in the configuration space of gravity? Hamiltonian constraints define an infinite-dimensional configuration space, superspace. Imposing geometric symmetries (such as homogeneity) on spacetime yields a finite-dimensional configuration space, mini-superspace. This reduction procedure can be viewed as a kind of fluid limit for (quantum) gravitational configurations. Generically, when many microscopic degrees of freedom are coarse-grained into a fluid description, symmetries like volume-preserving diffeomorphism invariance often emerge, independent of the details of the constituents. Therefore, it is expected that some common symmetry may emerge in various mini-superspaces, potentially carrying the universal symmetry of quantum gravity. Indeed, 2D Schrödinger symmetry was obtained in the homogeneous isotropic universe model and the spherically symmetric static vacuum model. In this paper, to investigate its robustness, we analyzed spherically symmetric static mini-superspace models coupled with matter fields at a classical level. Due to spherical symmetry and staticity, these become one-dimensional dynamical systems with Hamiltonian constraint $H=N\mathcal{H}\approx 0$, where the radial coordinate is the evolution parameter.
@@
First, it is crucial to investigate the robustness of Schrödinger symmetry in a broader context. Our current discussion is limited to cases reducible to the free-particle Hamiltonian without potential via canonical transformations. However, potentials of special forms, such as $V(x)=\frac{k}{x^{2}}$, also preserve Schrödinger symmetry Boyer:1974qt ; Nikitin:2017gtu ; Duval:2024eod . Therefore, it is important to consider such more general cases and investigate whether other matter fields, such as massive scalar field or Dirac field, possess Schrödinger symmetry, or reconsider whether the scalar-field model in terms of the $X$ coordinate in footnote 5 leads to the symmetry. Conversely, it would be interesting to assume such special potentials and examine the spacetime structure derived from them or consider the corresponding modified gravity theories.
@@
In particular, quantization of the scalar-field mini-superspace model is interesting. The interior region of the horizon in the Schwarzschild metric corresponds—via swapping two coordinates $t$ and rr— to a Kantowski-Sachs-type cosmological spacetime, where no matter field exists, of course. From this corresponding, quantum black-hole models are expected to exhibit a strong link with quantum cosmology models Modesto:2005zm . On the other hand, a scalar field can be used as a “clock” in relational formalism to quantize models Blyth:1975is ; Halliwell:1984eu ; Domagala:2010bm ; Ashtekar:2011ni . One of our interesting results in the scalar-field mini-superspace model is that, the JNW solution and the Kantowski-Sachs-type spacetime (with the causal-limit fluid) are unified into the single metric (V.3), meaning that the latter corresponds to the “interior” of the former. In particular, both possess the Schrödinger symmetry. Therefore, the quantization of this model based on the symmetry and this interior-exterior relationship should provide a promising avenue for connecting quantum black hole models and quantum universe models.
@@
Finally, it is crucial to examine the Schrödinger symmetry found in our mini-superspace models from the perspective of a hydrodynamic approximation derived from a full theory of quantum gravity. Ref.Oriti:2024elx focuses on the Schrödinger symmetry in the FLRW mini-superspace model and discusses the potential for removing cosmological singularities and realizing accelerated expansion through corrections to the quantum fluid equations that preserve the symmetryMarchetti:2025jze . It would be interesting to check whether a similar mechanism works in the spherically-symmetric static mini-superspace models. This test could serve as a crucial benchmark for assessing the universality of the hydrodynamic picture in quantum gravity. On the other hand, Ref.BenAchour:2023dgj pointed out the existence of Schrödinger symmetry in the spherically-symmetric static vacuum mini-superspace and suggested that nonlinear terms preserving the symmetry could serve as quantum corrections. However, the form of such corrections may depend on the choice of the lapse function. Indeed, the symmetry they identified corresponds to only one of the two possible choices of the lapse function and mini-superspace coordinate we found. It would be intriguing to consider different lapses and coordinates and investigate whether there are differences in the physical effects arising from the correction terms. Furthermore, we showed that Schrödinger symmetry persists even when adding matter fields. If the emergence of Schrödinger symmetry in such a mini-superspace is rigorously linked to the validity of the hydrodynamic approximation in a full quantum gravity theory incorporating matter fields, the findings of this study would provide significant implications.
@@ Acknowledgments
We would like to thank Jibril Ben Achour, Che-Yu Chen, Kristina Giesel, Philipp Höhn, Christy Kelly, Etera Livine, Francesco Sartini, and Shinji Tsujikawa for fruitful and inspiring discussions. T.S. is supported by Waseda University Special Research Projects (Nos. 2025C-493) and the 15th Early Bird Program of Waseda University. Y.Y. appreciated the hospitality of the members of the Institute for Theoretical Physics III at the Friedrich-Alexander-University Erlangen-Nuremberg, where he had a fruitful discussion and completed this work. Y.Y. was partially supported by Japan Society for the Promotion of Science (No.21K13929).
@@ Appendix A Mass and Electric Charge
where $g_{2d}$ is the determinant of the 2D metric normal to 2-sphere, and $dt_{\infty}=\frac{1}{k}dt$ is the time coordinate at $r=\infty$ with a constant $k$.
@@
Next, we show that in the Y-coordinate, the Komar formula (154) becomes the simpler formula (112). In terms of $dt_{\infty}=\frac{l_{s}\widetilde{\Pi}_{1}}{\bar{c}r_{0}}dt\equiv\frac{1}{k}dt$ and the lapse (86), the Y-metric (82) is expressed as
@@
Appendix B Lagrangian Interpretation of Symmetry under Hamiltonian constraints
In Sec.VI, we examined in detail the symmetries in the canonical formalism with Hamiltonian constraints, particularly providing an interpretation that certain symmetry transformations generate new theories. In this Appendix, we will study systems with Hamiltonian constraints in the action principle, explore the relationship between the canonical and Lagrangian descriptions, and show that the interpretation of generating new theories by symmetry transformations also holds in the action principles.
The mini-superspace models in this paper can generally be described by the following action:
$S[X,\mathcal{N}]=\int d\tau\mathcal{N}\left[\frac{1}{2\mathcal{N}^{2}}\mathcal{G}_{ab}(X)\frac{dX^{a}}{d\tau}\frac{dX^{b}}{d\tau}-\mathcal{V}(X)\right].$ (163)
Here, since the lapse $N(\tau)$ in the spacetime metric (26) is an arbitrary function of $\tau$ and may depend on $X^{a}(\tau)$, we have absorbed a function $f(X)$ that may appear alongside $N$, and introduced a new independent variable $\mathcal{N}(\tau)\equiv N(\tau)f(X(\tau))$. Indeed, say, the gravity action (28) and the scalar-field action in footnote 5 take this form. The variation of $\mathcal{N}$ yields
$\frac{\delta S}{\delta\mathcal{N}}=-\frac{1}{2\mathcal{N}^{2}}\mathcal{G}_{ab}\frac{dX^{a}}{d\tau}\frac{dX^{b}}{d\tau}-\mathcal{V}=0,$ (164)
which corresponds to the Hamiltonian constraint $\mathcal{H}\approx 0$. Therefore, the action (163) contains the effect of the Hamiltonian constraint. On the other hand, the variation of $X^{a}$ provides the equation of motion for $X^{a}$ (geodesic equation in the mini-superspace), which is equivalent to the canonical equation for the Hamiltonian $H=\mathcal{N}\mathcal{H}$. Solving this equation together with the constraint (164) determines a physical solution $X^{a}(\tau)$. Therefore, all the equations are derived from the action principle of the non gauge-fixed action (163), which determines the full dynamics and defines the theory. In this sense, we call it the full action.
Alternatively, we can fix the lapse as $\mathcal{N}(X)=1$ at the action level and consider the gauge-fixed action:
$S_{\mathrm{fix}}[X]=\int d\tilde{\tau}\left(\frac{1}{2}\mathcal{G}_{ab}(X)\frac{dX^{a}}{d\tilde{\tau}}\frac{dX^{b}}{d\tilde{\tau}}-\mathcal{V}(X)\right),$ (165)
where $\tilde{\tau}$ denotes the time coordinate in this gauge. In the action principle of this gauge-fixed action, the Hamiltonian constraint (164) cannot be derived while the $X$ equation of motion can in this gauge. Therefore, considering only the gauge-fixed action $S_{\rm fix}$ corresponds to analyzing the $X$ equation of motion, equivalent to the canonical equations for $H=\mathcal{H}$, without imposing $\mathcal{H}\approx 0$. In order to obtain a physical solution, we need to impose the constraint (164) (for $\mathcal{N}=1$) by hand and solve it with the $X$ equations.
Now, we examine the transformations generated by Noether charges in the full action and the gauge-fixed action. In this paper (the canonical formalism), Noether charges are defined by the condition (6), which leaves the gauge-fixed action invariant. Indeed, this condition guarantees the invariance of $H$ for such transformations (see Eq.(120)), and the canonical form of the gauge-fixed action $S_{\rm fix}^{\rm can}[X,P]=\int d\tilde{\tau}(P_{a}\dot{X}^{a}-H)$ is invariant, since for general canonical transformations, the symplectic structure $P_{a}dX^{a}$ is invariant. However, this does not mean the invariance of the full action $S[X,\mathcal{N}]$. This is natural, since the gauge-fixed action does not have the effect of the Hamiltonian constraint. In fact, not all such transformations leave the full action invariant (see below). Therefore, some transformations generated by Noether charges satisfying (6) can change the theory.
To illustrate these points concretely, let us consider a simple case where $\mathcal{G}_{ab}$ is a constant invertible matrix and $\mathcal{V}(X)={\rm const.}=\lambda$:
$S[X,\mathcal{N}]=\int d\tau\mathcal{N}\left(\frac{1}{2\mathcal{N}^{2}}\mathcal{G}_{ab}\frac{dX^{a}}{d\tau}\frac{dX^{b}}{d\tau}-\lambda\right).$ (166)
If we fix $\mathcal{N}=1$, this becomes
$S_{\mathrm{fix}}[X]=\int d\tilde{\tau}\left(\frac{1}{2}\mathcal{G}_{ab}\frac{dX^{a}}{d\tilde{\tau}}\frac{dX^{b}}{d\tilde{\tau}}-\lambda\right).$ (167)
Note that this form of the action is derived from the Hamiltonian (118) via the Legendre transformation.
As an example, we consider the Galilean boost transformation given by
$X^{a}\rightarrow X^{\prime a}=X^{a}-v^{a}\eta,$ (168)
where $v^{a}$ is constant and $\eta=\eta(\tau):=\int^{\tau}d\tau^{\prime}\mathcal{N}(\tau^{\prime})$. For the gauge fixing $\mathcal{N}=1$, this becomes
$X^{a}\rightarrow X^{\prime a}=X^{a}-v^{a}\tilde{\tau}$. (169)
Under this transformation, the gauge-fixed action (167) transforms as $S_{\mathrm{fix}}\to S^{\prime}_{\mathrm{fix}}=S_{\mathrm{fix}}[X^{\prime}]$:
$S^{\prime}_{\mathrm{fix}}$ $=\int d\tilde{\tau}\left(\frac{1}{2}\mathcal{G}_{ab}\left(\frac{dX^{a}}{d\tilde{\tau}}-v^{a}\right)\left(\frac{dX^{b}}{d\tilde{\tau}}-v^{b}\right)-\lambda\right)$
$=\int d\tilde{\tau}\Bigg(\frac{1}{2}\mathcal{G}_{ab}\frac{dX^{a}}{d\tilde{\tau}}\frac{dX^{b}}{d\tilde{\tau}}-\lambda$
$\;\;\;\;\;\;\;\;\;\;\;\;+\frac{d}{d\tilde{\tau}}\left(-\mathcal{G}_{ab}v^{a}X^{b}\right)+\frac{1}{2}\mathcal{G}_{ab}v^{a}v^{b}\Bigg)$
$=S_{\rm fix}+{\rm(surface~terms)}$ (170)
[…]
On the other hand, the full action (166) changes for the boost (168) as $S\to S^{\prime}=S[X^{\prime},\mathcal{N}]$:
$S^{\prime}$ $=\int d\tau\mathcal{N}\left(\frac{1}{2\mathcal{N}^{2}}\mathcal{G}_{ab}\left(\frac{dX^{a}}{d\tau}-v^{a}\mathcal{N}\right)\left(\frac{dX^{b}}{d\tau}-v^{b}\mathcal{N}\right)-\lambda\right)$
$=\int d\tau\Bigg(\mathcal{N}\left[\frac{1}{2\mathcal{N}^{2}}\mathcal{G}_{ab}\frac{dX^{a}}{d\tau}\frac{dX^{b}}{d\tau}-\lambda+\frac{1}{2}\mathcal{G}_{ab}v^{a}v^{b}\right]$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~-\frac{d}{d\tau}\left(\mathcal{G}_{ab}v^{a}X^{b}\right)\Bigg)$
$=\int d\tau\mathcal{N}\left[\frac{1}{2\mathcal{N}^{2}}\mathcal{G}_{ab}\frac{dX^{a}}{d\tau}\frac{dX^{b}}{d\tau}-\lambda+\frac{1}{2}\mathcal{G}_{ab}v^{a}v^{b}\right]$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+{\rm(surface~term)}.$ (171)
Unlike the gauge-fixed case, the term $\frac{1}{2}\mathcal{N}\mathcal{G}_{ab}v^{a}v^{b}$ cannot be ignored due to the presence of $\mathcal{N}$. Consequently, the Hamiltonian constraint is modified as
$\frac{\delta S^{\prime}}{\delta\mathcal{N}}=-\frac{1}{2\mathcal{N}^{2}}\mathcal{G}_{ab}\frac{dX^{a}}{d\tau}\frac{dX^{b}}{d\tau}-\lambda+\frac{1}{2}\mathcal{G}_{ab}v^{a}v^{b}=0,$ (172)
which leads to the new Hamiltonian constraint (130). Thus, this symmetry transformation changes the full action, modifying the constraint. Note that the modification is just a constant shift, $\lambda\to\lambda-\frac{1}{2}\mathcal{G}_{ab}v^{a}v^{b}$, which does not affect the variation of $X^{a}$. This corresponds to the fact that the time evolution operator $H_{0}$ remains invariant in the canonical formalism (see Sec.VI.1).
Note that there exists another interpretation of such transformations: if $\mathcal{N}$ is also transformed properly, the full action can be invariant, leading to dynamical symmetries. Indeed, this idea works for the generators $Q_{+}$, $Q_{0}$, and $Q_{-}$ corresponding to conformal transformations Achour2019 . For the Galilean boost $B^{a}$, however, to the best of our knowledge, there exists no such interpretation.
Thus, the interpretation of generating new theories by symmetry transformations are consistent both in the canonical and Lagrangian formalisms. We here stress that the advantage of the canonical-formalism analysis in Sec.VI lies in its detailed examination of of the subtle points associated with gauge constraints (point I, II, III).